Square root property to solve the equation:
(1-3x)^2=23

Answer by stanbon(75887)

Square root property to solve the equation:
(1-3x)^2=23
-------------------
1-3x = sqrt(23) or 1-3x = -sqrt(23)
-----
3x = 1-sqrt(23) or 3x = 1+sqrt(23)
------------------
x = (1-sqrt(23))/3 or x = (1+sqrt(23))/3
======================
Cheers,
Stan H.